# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ Abstract.
In a previous work, we investigated the analytic continuation of the bulk Poisson operator of Liouville conformal field theory on the holomorphic part of Fock space and used it to construct irreducible representations of the Virasoro algebra at the degenerate values of the conformal weights. Here, we study two cases where the Poisson operator admits some simple poles on the Kac table: the bulk Poisson operator on the full Fock space (both holomorphic and anti-holomorphic part), and the boundary Poisson operator on the Fock space. As a consequence, the derivative of top singular vector does not vanish, and we can identify it with a scalar multiple of a primary field of same conformal weight. These are known as higher-equations of motions in physics and have been studied in connection with minimal gravity.
@@ 1.1. Motivation and background
In Segal’s axiomatisation [Seg04], one basic ingredient of a conformal field theory (CFT) is a Hilbert space $\mathcal{H}$ together with a (projective) representation of the “semigroup of annuli”, the space of complex annuli with analytically parametrised boundaries. Provided they exist, the generators of this semigroup should form a family of (unbounded) operators on $\mathcal{H}$ representing the Virasoro algebra (in a suitable sense). In most cases, one asks for two commuting representations $(\mathbf{L}_{n},\widetilde{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$, one for each “sector” (holomorphic and antiholomorphic). The Hamiltonian of the theory is then the zero mode $\mathbf{H}=\mathbf{L}_{0}+\widetilde{\mathbf{L}}_{0}$, and it is self-adjoint for a unitary theory. Diagonalising this Hamiltonian allows one to write $\mathcal{H}$ as a direct sum of highest-weight representations, which are called the spectrum of the CFT. In rational CFTs, the spectrum is discrete, and all the representations are degenerate, in the sense that the highest-weights lie in the Kac table. Moreover, it is often taken as an axiom that the representations in the spectrum are irreducible, implying that all singular vectors vanish.
The case of Liouville CFT is different since there is a continuous spectrum of representations, none of which being degenerate. The theory depends on a complex parameter $\gamma$, and a probabilistic construction is available for $\gamma\in(0,2)$ [DKRV16, DRV16, GRV19], which is the framework of this paper. The diagonalisation of the Hamiltonian was done in [GKRV23], and the representations were constructed in [${}^{+}$23]. The states in the spectrum are labelled $(\Psi_{Q+ip,\nu,\widetilde{\nu}})_{p\in\mathbb{R}_{+},\nu,\widetilde{\nu}\in\mathcal{T}}$, where $\mathcal{T}$ denotes the set of integer partitions (a.k.a. Young diagrammes), and $Q=\frac{2}{\gamma}+\frac{\gamma}{2}$. These states admit an analytic continuation to all complex values of $Q+ip=\alpha$ [${}^{+}$23]. Away from the Kac table $kac^{-}=(1-\mathbb{N}^{*})\frac{\gamma}{2}+(1-\mathbb{N}^{*})\frac{2}{\gamma}$, the linear span of $(\Psi_{\alpha,\nu,\widetilde{\nu}})_{\nu,\widetilde{\nu}\in\mathcal{T}}$ is isomorphic to a tensor product of Verma modules $\mathcal{V}_{\alpha}\otimes\bar{\mathcal{V}}_{\alpha}$: namely, the states $\Psi_{\alpha,\nu,\widetilde{\nu}}$ are the Virasoro descendants of the primary state $\Psi_{\alpha}=\Psi_{\alpha,\eset,\eset}$. For the degenerate values of the conformal weight, the linear span of $\Psi_{\alpha,\nu,\eset}$ forms a degenerate representation which is isomorphic to the irreducible quotient of the Verma module by the maximal proper submodule [BW23]. This is the algebraic property that is required for the null-vector equations to hold, which are key to the solutions of many CFTs [BPZ84, Tes01, KRV19, KRV20].
Our earlier work addressed the case where $\widetilde{\nu}=\eset$ [BW23], namely for states which are primaries with respect to the antiholomorphic representation $\widetilde{\mathbf{L}}_{n}$. The proof relies on the analytic continuation (to the $\alpha$-plane) of the “Poisson operator” $\mathcal{P}_{\alpha}$, and the observation that the vanishing of singular vectors is equivalent to the regularity of this operator. Informally, $\mathcal{P}_{\alpha}$ maps the free field eigenstates (where everything is known) to the Liouville ones. The case where $\widetilde{\nu}\neq\eset$ is more complicated and is the subject of conjectures in the physics literature [Zam04], which we summarise now. For each $\alpha_{r,s}=(1-r)\frac{\gamma}{2}+(1-s)\frac{2}{\gamma}\in kac^{-}$, there exists a linear combination of Virasoro generators $\mathbf{S}_{\alpha_{r,s}}$ at level $rs$ such that $\mathbf{S}_{\alpha_{r,s}}\Psi_{\alpha_{r,s}}=0$ [BW23]. It is customary to normalise this linear combination such that the coefficient of $\mathbf{L}_{-1}^{rs}$ equals 1. In fact, the first derivative is non-zero at $\alpha_{r,s}$, so that $\mathbf{S}_{\alpha_{r,s}}\Psi_{\alpha_{r,s}}^{\prime}\neq 0$, where the prime indicates derivative in $\alpha$ (recall $\Psi_{\alpha}$ is analytic in $\alpha$). According to [Zam04, Equation (4.5)],
$\widetilde{\mathbf{S}}_{\alpha_{r,s}}\mathbf{S}_{\alpha_{r,s}}\Psi_{\alpha_{r,s}}^{\prime}=B_{r,s}\Psi_{\alpha_{-r,s}},$ (1.1)
for some non-zero scalar $B_{r,s}$, for which there exists an exact formula [Zam04, Equation (5.17)]. We emphasise that $\alpha_{-r,s}$ does not belong to the Kac table, so $\Psi_{\alpha_{-r,s}}$ is not a degenerate state. This identity is applied by Zamolodchikov in the context of minimal gravity in order to reduce some integrals of correlation functions over the moduli space of curves to boundary integrals on the compactification divisor [Zam04, BZ06]. The relation (1.1) is called the higher-equations of motion (HEM), and this paper studies the validity of this formula at level 2: $(r,s)\in\{(1,2),(2,1)\}$.
In our previous work, we expressed the Virasoro descendants of the state $\Psi_{\alpha}$ as (a linear combination of) screening integrals with respect to singular weights, which we dubbed singular integrals. These integrals are of the form $\mathcal{I}_{r,\boldsymbol{s},\widetilde{\boldsymbol{s}}}(\alpha)=\int_{\mathbb{D}^{r}}\Psi_{\alpha}(w_{1},...,w_{r})\prod_{j=1}^{r}\frac{|\mathrm{d}w_{j}|^{2}}{w_{j}^{s_{j}}\bar{w}_{j}^{\widetilde{s}_{j}}}$, with the notation explained in (2.1). The integrals we studied in [BW23] all have $\widetilde{\boldsymbol{s}}=(0,...,0)$, and they extend analytically to the whole $\alpha$-plane. When $\widetilde{\boldsymbol{s}}$ is non-trivial, these integrals may develop some poles on the Kac table. Thus, the integrals $\mathcal{I}_{r,\boldsymbol{s},\widetilde{\boldsymbol{s}}}(\alpha)$ are really computing a series expansion of $\Psi_{\alpha}(w_{1},...,w_{r})$ in the neighbourhood of 0, when $\alpha$ belongs to the Kac table. The analyticity of these integrals away from the Kac table is a manifestation of the fact that $\Psi_{\alpha}(w_{1},...,w_{r})$ has a fractional behaviour for generic $\alpha$. The study of the HEMs is therefore of independent mathematical interest as it gives non-trivial information on the local structure of Gaussian multiplicative chaos with degenerate insertions.
@@ 1.2. Main results
where *$\mathbb{N}^{*}=\mathbb{Z}_{>0}$. Note that $kac^{+}=2Q-kac^{-}$, and $kac^{-}\subset(-\infty,0]$. We use the notation
$\alpha_{r,s}:=(1-r)\frac{\gamma}{2}+(1-s)\frac{2}{\gamma},$
for all $r,s\in\mathbb{Z}$. In practice, we will only focus on the case $(r,s)\in\{(1,2),(2,1)\}$. For all $\alpha\in\mathbb{C}$, we define:
$\Delta_{\alpha}:=\frac{\alpha}{2}\left(Q-\frac{\alpha}{2}\right).$
@@ 1.2.1. Bulk HEM
Our first result is to etablish the HEMs (1.1) at level 2 (i.e. $(r,s)\in\{(1,2),(2,1)\}$). For the next statement, recall that $\Psi_{\alpha}$ is a primary field of weight $\Delta_{\alpha}$ which is analytic in $\alpha$ and belongs to the weighted space $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ for $\beta>|Q-\mathrm{Re}(\alpha)|$. The operators $(\mathbf{L}_{n})_{n\in\mathbb{Z}}$ can be defined as bounded operators $e^{-\beta c}\mathcal{D}(\mathcal{Q})\to e^{-\beta c}\mathcal{D}^{\prime}(\mathcal{Q})$ and form a Virasoro representations. These constructions are detailed in Section 2.1.
@@
Theorem 1.1 confirms the prediction of Zamolodchikov, except the $(2,1)$-equation in the regime $\gamma\in(\sqrt{2},2)$ (where $\alpha_{-2,1}>Q$). The reason is a freezing phenomenon which makes the equation crash down (see Open question 3). In the other cases, the scalar multiple of the primary field coincides with [Zam04, Equation (5.17)] [^1] With the dictionary $(r,s)\leftrightarrow(n,m)$ and $\frac{\gamma}{2}\leftrightarrow b$, and [Zam04] uses the notation $\gamma(z)=\frac{\Gamma(z)}{\Gamma(1-z)}$.. Equation (1.3) has a non-trivial limit as $\gamma\to\sqrt{2}$ from below. We expect that the $(2,1)$-equation holds in this case, but it requires a more detailed analysis of Gaussian multilpicative chaos (see Open question 1).
We will prove Theorem 1.1 in Section 2. As in [BW23], the strategy is to study the analytic continuation of the Poisson operator. The main difference is that the extension has a simple pole at $\alpha_{1,2}$ or $\alpha_{2,1}$. The RHS of (1.3) is then identified with the residue of the Poisson operator. At the time of writing, we are aware of an online manuscript studying the $(2,1)$-HEM for correlation functions on the Riemann sphere [AR]. Their proof is close in spirit to the existing literature on BPZ equations in Liouville CFT, while draws from techniques introduced in [BW23]. Moreover, the scalar multiple of the primary field of [AR] differs from ours (and Zamolodchikov’s); this seems to be due to the omission of a pole in the study of the correlation function of the singular state.
[…]
$(\alpha_{0}^{2}\mathcal{L}_{-2}+\mathcal{L}_{-1}^{2})\mathcal{B}(\otimes_{j=0}^{m}\Psi_{\alpha_{j}})=0,$
[…]
$\mathcal{S}_{\alpha}:=\alpha^{2}\mathcal{L}_{-2}+\mathcal{L}_{-1}^{2};\qquad\widetilde{\mathcal{S}}_{\alpha}:=\alpha^{2}\widetilde{\mathcal{L}}_{-2}+\widetilde{\mathcal{L}}_{-1}^{2}.$
Corollary 1.2.
Let $(\alpha_{0},...,\alpha_{m})\in\mathbb{C}^{m+1}$ satisfying the Seiberg bounds, with $\alpha_{0}\in\{\alpha_{1,2},\alpha_{2,1}\}$, and set $\boldsymbol{\Psi}:=\otimes_{j=1}^{m}\Psi_{\alpha_{j}}$. Let $\mathcal{B}:\mathcal{T}^{\epsilon}_{\Sigma,m+1,\uparrow}\to\mathcal{D}^{\prime}(\mathcal{Q})^{\otimes(m+1)}$ be a conformal block in the sense of [BW23, Section 4].
  1. 1.
    We have
    $\widetilde{\mathcal{S}}_{\alpha_{1,2}}\mathcal{S}_{\alpha_{1,2}}\mathcal{B}(\Psi_{\alpha_{1,2}}^{\prime}\otimes_{j=1}^{m}\Psi_{\alpha_{j}})=\pi\mu\frac{8}{\gamma^{3}}\left(1-\frac{\gamma^{2}}{4}\right)^{2}\mathcal{B}(\Psi_{\alpha_{-1,2}}\otimes_{j=1}^{m}\Psi_{\alpha_{j}}).$
  2. 2.
    We have
    $\widetilde{\mathcal{S}}_{\alpha_{2,1}}\mathcal{S}_{\alpha_{2,1}}\mathcal{B}(\Psi_{\alpha_{2,1}}^{\prime}\otimes\boldsymbol{\Psi})=\left\{\begin{aligned} &-\frac{\gamma^{5}}{32}\left(\pi\mu\frac{\Gamma(\frac{\gamma^{2}}{4})}{\Gamma(1-\frac{\gamma^{2}}{4})}\right)^{2}\frac{\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(\frac{\gamma^{2}}{2})}\mathcal{B}(\Psi_{\alpha_{-2,1}}\otimes\boldsymbol{\Psi}),&\text{if }\gamma<\sqrt{2};\\ &0,&\text{if }\gamma>\sqrt{2}.\end{aligned}\right.$
@@ 1.2.2. Boundary HEMs
Liouville CFT has a boundary counterpart, which has been under intense investigation in the probability community over the past few years. On the one hand, there is a vast integrability programme which culminated in the derivation of all structure constants of the theory [ARSZ23]. These formulas are the analogues of the celebrated DOZZ formula of bulk LCFT [KRV20]. On the other hand, there is work in progress studying the conformal bootstrap for boundary LCFT [GRVW]. Our results could be formulated compactly (as in Theorem 1.1) using the technology of [GRVW], but we will instead follow a more pedestrian route until this work appears online.
The details of boundary LCFT are recalled in Section 3.1. In a nutshell, the states are associated to the half-disc $\mathbb{D}\cap\mathbb{H}$. The interior of this half-disc bears a (bulk) cosmological cosmological constant $\mu>0$. Additionally, there are two boundary cosmological constants $\mu_{\mathrm{L}},\mu_{\mathrm{R}}>0$ associated to the “left” interval $\mathbb{I}^{-}=(-1,0)$ and the “right” interval $\mathbb{I}^{+}=(0,1)$ respectively. There is a family of free field representatiosn $(\mathbf{L}_{n}^{0,\alpha})_{n\in\mathbb{Z}}$ (the Sugawara construction), and we define $\mathbf{S}^{0}_{\alpha}:=\alpha^{2}\mathbf{L}_{-2}^{0,\alpha}+(\mathbf{L}_{-1}^{0,\alpha})^{2}$.
We refer to Section 3.1 for precise definitions of the objects appearing in the statement of the next theorem. Briefly, $\Psi_{\alpha}^{\partial}$ is a state which is analytic for $\alpha$ in a neighbourhood of $(-\infty,Q)$, as an element of a weighted Hilbert space $e^{-\beta c}\mathcal{H}$. This state is expected to be a highest-weight state for a Virasoro representation $(\mathbf{L}_{n})_{n\in\mathbb{Z}}$ whose construction should be similar to the bulk one. The evaluation of the expression $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ (see (3.2) for the definition) at the degenerate weights $\alpha_{1,2}$, $\alpha_{2,1}$, should correspond to the value of the (level 2) Liouville singular vector in this representation. In other words, we expect that $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})=(\alpha^{2}\mathbf{L}_{-2}+\mathbf{L}_{-1}^{2})\Psi_{\alpha}^{\partial}$ for some Virasoro representation $(\mathbf{L}_{n})_{n\in\mathbb{Z}}$ which is yet to bet constructed. Combining with the results of [GRVW], the next theorem will give us the value of the singular vectors at level 2. We see that this singular vector is always proportional to a primary field of different weight, as in the bulk case (with the constant possibly equal to zero).
@@
The expression $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})\in e^{-\beta c}\mathcal{H}$ admits an analytic extension which is regular in the neighbourhood of $\alpha_{1,2}$ and $\alpha_{2,1}$. Moreover,
@@
    $\Phi_{\alpha_{2,1}}(\mathbf{S}_{2Q-\alpha_{2,1}}^{0}\mathds{1})=\left\{\begin{aligned} &\frac{\gamma^{3}}{8}\left(\mu_{\mathrm{L}}^{2}-2\mu_{\mathrm{L}}\mu_{\mathrm{R}}\cos(\pi\frac{\gamma^{2}}{4})+\mu_{\mathrm{R}}^{2}-\mu\sin(\pi\frac{\gamma^{2}}{4})\right)\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}\Psi_{\alpha_{-2,1}}^{\partial}&\text{ if }\gamma<\sqrt{2};\\ &0&\text{if }\gamma>\sqrt{2}.\end{aligned}\right.$ (1.5)

### v2

@@ Abstract.
We prove conjectures of Zamolodchikov and Belavin–Belavin in Liouville conformal field theory (CFT), which are generalisations of the celebrated Belavin–Polyakov–Zamolodchikov equations known as the higher equations of motion. Algebraically, these equations give examples of non-zero singular states in Virasoro modules, which is a relatively rare phenomenon in the physical study of CFT. In probability theory, these equations and their variants have been instrumental in the rigorous derivation of the structure constants of Liouville CFT in the unit disc.
The proof builds on a previous work of ours studying the analytic continuation of the Poisson operator of Liouville theory. The main novelty is that this operator admits poles on the Kac table, and the higher equations of motions are obtained via a residue computation.
Key words and phrases:
Gaussian free field, Gaussian multiplicative chaos, Virasoro algebra, conformal field theory.
2020 Mathematics Subject Classification:
60G60; 17B68; 81T40
@@ 1.1. Motivation and background
In the last decade, a series of breakthrough results led to a rigorous construction of Liouville conformal field theory (CFT), using probabilistic techniques based on the Gaussian free field and Gaussian multiplicative chaos [DKR+16a, HRV18, DRV16b, GRV19]. Prior to this, the focus in physics was put on the algebraic structure of Liouville CFT (see e.g. [TES01, NAK04, RIB14] for reviews in physics) rather than its analytic aspects (with some notable exceptions exploring the connections between Liouville theory and the complex-analytic formalism of CFT proposed by Friedan–Shenker [FS87], see e.g. [TT06] and references therein). Based on the assumption of conformal symmetry and using results from representation theory, the algebraic approach has striking predictions (such as exact formulas for certain correlations and the “conformal bootstrap hypothesis” [DO94, ZZ96, TES03, BPZ84]) which were verified in the probabilistic setting [KRV20, GKR+24a, GKR+21]. The cornerstone of this integrable structure is a family of PDEs satisfied by correlation functions involving so-called degenerate fields: these are the celebrated BPZ equations [BPZ84], proved in [KRV19] for the Liouville CFT. The existence of such PDEs is a consequence of conformal symmetry (in the form of Ward identities) and the physically motivated assumption that “states with vanishing norm must vanish identically” (see [BW23, Section 1.1] for a discussion of the meaning of this phrase). Still, the BPZ equations are down-the-road consequences of conformal symmetry, and more recent works in probability theory have focused on the construction of the whole algebraic structure expected for a CFT: a family of (unbounded) operators acting on the Hilbert space of the theory, which represents two commuting copies of the Virasoro algebra and exponentiate to a projective representation of Segal’s semigroup of annuli [GKR+24a, BGK+24a, BGK+24b]. The existence of such a structure constrains the Liouville Hilbert space through representation theory [GKR+24a, BGK+24a], and the BPZ equations can be reformulated as a statement on the algebraic structure of certain representations. While this structure is usually taken as an axiom in physics, it needs to be proved in the probabilistic representation [BW23, Theorem 1.2].
This paper can be seen as a continuation of [BW23], and is concerned with two cases where this basic axiom does not hold: these are known as the higher-equations of motion (HEM) and have been predicted in [ZAM04, BB10]. Aside from the intrinsic interest of these equations (which have applications in minimal gravity and intersection theory [BZ06, AB22, AC24]), they stress the importance of not taking the BPZ equations for granted. In the probabilistic literature, some hints for the HEMs appeared in [ANG23], where the (level 2) BPZ equations for Liouville CFT on the disc are proved under some conditions on the cosmological constants. Relaxing this condition, Cerclé proves the full HEMs for correlation functions in the disc, in an independent work which appeared online a few weeks after the present one. The (level 2) HEMs as stated in [BB10] (i.e. at the level of states rather than correlations) is the topic of Theorem 1.2 below. As noted by both Cerclé and us, these HEMs do not hold for some range of the parameters (second item of Theorem 1.2, case $\gamma>\sqrt{2}$). The special case proved by Ang has been instrumental in a wide programme aiming at the computation of all structure constants of boundary Liouville CFT [ARS+23], following predictions from physics [FZZ00, PT02, HOS01]. In addition to the boundary theory, Theorem 1.1 studies the bulk version of these HEMs, which were predicted earlier than the boundary versions by Zamolodchikov [ZAM04] (and are not addressed in [ANG23, CER25]). Finally, we mention that Cerclé–Huguenin recently announced a version of the HEMs for the boundary $\mathfrak{sl}_{3}$-Toda CFT [CH25].
In the remainder of this introduction, we state our main results and provide an outline of the proofs, as well as some open questions. The proofs of Theorem 1.1 and Theorem 1.2 follow the same strategy and can be read independently (with quite some extra technicalities for the bulk version). In particular, the reader only interested in the boundary HEM may skip directly to Section 3 only.
@@ 1.2. Main results
where $\mathbb{N}^{*}=\mathbb{Z}_{>0}$. Note that $kac^{+}=2Q-kac^{-}$, and $kac^{-}\subset(-\infty,0]$. We use the notation $\alpha_{r,s}:=(1-r)\frac{\gamma}{2}+(1-s)\frac{2}{\gamma}$ for all $r,s\in\mathbb{Z}$, and note that:
$\alpha_{1,2}=-\frac{2}{\gamma}\qquad\text{and}\qquad\alpha_{2,1}=-\frac{\gamma}{2}.$
For all $\alpha\in\mathbb{C}$, we define the conformal weight
$\Delta_{\alpha}:=\frac{\alpha}{2}\left(Q-\frac{\alpha}{2}\right).$ (1.1)
@@ 1.2.1. Bulk HEM
In order to state our main result, we need to introduce some notions of Liouville CFT, which are recalled in greater detail in Section 2.1.
The Hilbert space of Liouville CFT $\mathcal{H}$ is the $L^{2}$-space of a Gaussian measure with values in the space of distributions on the unit circle. It comes with a Hilbert space $\mathcal{H}$ and a family of densely-defined, unbounded operators $(\mathbf{L}_{n},\widetilde{\mathbf{L}}_{n})_{n\in\mathbb{Z}}$ on $\mathcal{H}$, which represent two commuting families of the Virasoro algebra in a suitable sense. The Hamiltonian is the operator $\mathbf{H}=\mathbf{L}_{0}+\widetilde{\mathbf{L}}_{0}$, which is self-adjoint and induces a positive definite quadratic form with dense domain $\mathcal{D}(\mathcal{Q})\subset\mathcal{H}$. Then, for all $\beta>0$ and all $n\in\mathbb{Z}$, the Virasoro generators are bounded as operators $\mathbf{L}_{n},\widetilde{\mathbf{L}}_{n}:e^{-\beta c}\mathcal{D}(\mathcal{Q})\to e^{-\beta c}\mathcal{D}^{\prime}(\mathcal{Q})$, where $\mathcal{D}^{\prime}(\mathcal{Q})$ is the continuous dual to $\mathcal{D}(\mathcal{Q})$ (see Section 2.1 for a precise definition of the weighting). Furthermore, for each pair of Young diagrams $\nu,\widetilde{\nu}$ (see Section 2.1 for the definition), there exists an analytic family of generalised eigenstates $\alpha\mapsto\Psi_{\alpha,\nu,\widetilde{\nu}}\in e^{-\beta c}\mathcal{D}(\mathcal{Q})$ (with $\beta>|Q-\mathrm{Re}(\alpha)|$), such that $\Psi_{\alpha,\nu,\widetilde{\nu}}$ is obtained by acting on $\Psi_{\alpha}$ with the Virasoro generators. In particular, these operators can be viewed as endomorphisms of the vector space $\mathcal{W}_{\alpha}$ (algebraically) spanned by the generalised eigenstates, which is the point of view adopted in the next statement (see Section 2.1 and [BGK+24a, Section 4.4] for some details on these properties). Since $\Psi_{\alpha}$ is analytic with values in $e^{-\beta c}\mathcal{D}(\mathcal{Q})$ for some $\beta$, we can consider its complex derivative $\Psi_{\alpha}^{\prime}:=\partial_{\alpha}\Psi_{\alpha}\in e^{-\beta c}\mathcal{D}(\mathcal{Q})$. In the next statement, $\beta$ is any fixed number satisfying the above property.
@@
Theorem 1.1 confirms the prediction of Zamolodchikov, except the $(2,1)$-equation in the regime $\gamma\in(\sqrt{2},2)$ (where $\alpha_{-2,1}>Q$). The reason is a freezing phenomenon which makes the equation break down (see Open question 3). In the other cases, the scalar multiple of the primary field coincides with [ZAM04, Equation (5.17)] [^1] With the dictionary $(r,s)\leftrightarrow(n,m)$ and $\frac{\gamma}{2}\leftrightarrow b$, and [ZAM04] uses the notation $\gamma(z)=\frac{\Gamma(z)}{\Gamma(1-z)}$.. Equation (1.3) has a non-trivial limit as $\gamma\to\sqrt{2}$ from below. We expect that the $(2,1)$-equation holds in this case, but it requires a more detailed analysis of Gaussian multiplicative chaos (see Open question 1).
We will prove Theorem 1.1 in Section 2. As in [BW23], the strategy is to study the analytic continuation of the Poisson operator. The main difference is that the extension has a simple pole at $\alpha_{1,2}$ or $\alpha_{2,1}$. The RHS of (1.3) is then identified with the residue of the Poisson operator. At the time of writing, we are aware of an online manuscript studying the $(2,1)$-HEM for correlation functions on the Riemann sphere [1]. Their proof is close in spirit to Cerclé’s proof of the boundary HEM, but the scalar multiple of the primary field of [1] differs from ours (and Zamolodchikov’s); this seems to be due to the omission of a pole in the study of the correlation function of the singular state.
Theorem 1.1 is an equality of states in Liouville CFT. The link between states and correlation functions in Liouville CFT is given by the so-called amplitudes introduced in [GKR+21]: a correlation function on an arbitrary compact Riemann surface can be obtained as the inner-product of a state with the amplitude of a surface with boundary. On the other hand, [BGK+24a, BGK+24b] relate the action of the Virasoro algebra to deformations of the complex structure. In a work in progress of the first author and Guillarmou, Kupiainen, Rhodes [8], it is shown in full generality how to relate the Virasoro algebra to deformations of the complex structure on an arbitrary compact Riemann surface. Concretely, the machinery developed in [8] implies that conformal blocks involving Virasoro descendants are obtained from conformal blocks of primary fields by applying a partial differential operator on the Teichmüller space of the surface. In particular, Theorem 1.1 implies that conformal blocks involving the state $\mathbf{S}_{\alpha_{r,s}}\widetilde{\mathbf{S}}_{\alpha_{r,s}}\Psi_{\alpha_{r,s}}^{\prime}$ (for $rs=2$) satisfy a certain PDE on Teichmüller space. This generalises the celebrated BPZ equations. Since this relies on the work in progress [8], we refrain from writing a full statement for the time being.
@@ 1.2.2. Boundary HEMs
Liouville CFT has a boundary counterpart, which has been under intense investigation in the probability community over the past few years. On the one hand, there is a vast integrability programme which culminated in the derivation of all structure constants of the theory [ARS+23]. These formulas are the analogues of the celebrated DOZZ formula of bulk LCFT [KRV20]. On the other hand, there is work in progress studying the conformal bootstrap for boundary LCFT [GRW24b]. Our results could be formulated compactly (as in Theorem 1.1) using the technology of a work still in progress, but we will instead follow a more pedestrian route until this work appears online.
The details of boundary LCFT are recalled in Section 3.1. The Hilbert space $\mathcal{H}$ is the $L^{2}$-space of a Gaussian measure with values in the space of distributions on the half-circle $\mathbb{S}^{1}\cap\mathbb{H}$. In addition to the bulk cosmological constant $\mu$, it depends on two additional constants $\mu_{\mathrm{L}},\mu_{\mathrm{R}}>0$ associated with the intervals $\mathbb{I}^{-}=(-1,0)$ and $\mathbb{I}^{+}=(0,1)$ respectively. The Sugawara construction $(\mathbf{L}_{n}^{0,\alpha})_{n\in\mathbb{Z}}$ is a family of densely defined, unbounded operators which represent the Virasoro algebra. We define $\mathbf{S}^{0}_{\alpha}:=\alpha^{2}\mathbf{L}_{-2}^{0,\alpha}+(\mathbf{L}_{-1}^{0,\alpha})^{2}$.
As in the bulk theory, one can define a family of states $\alpha\mapsto\Psi_{\alpha}^{\partial}\in e^{-\beta c}\mathcal{H}$ depending analytically on $\alpha$ and with values in the weighted Hilbert space $e^{-\beta c}\mathcal{H}$ with $\beta>|Q-\mathrm{Re}(\alpha)|$ (see Section 3.1 for the precise definition). In Section 3.1, we introduce a map which informally speaking sends the “free field state” $\mathbf{S}^{0}_{\alpha}\mathds{1}$ to the “Liouville state” $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})\in e^{-\beta c}\mathcal{H}$. Analogously to [BGK+24a, Theorem 4.5], it will be shown in a work in progress [33] that this state is the descendant of a primary field $\Psi_{\alpha}^{\partial}$ for some Virasoro representation $(\mathbf{L}_{n})_{n\in\mathbb{Z}}$. In the meantime, we introduce this state $\Psi_{\alpha}^{\partial}$ in Section 3.1, and show its analyticity for $\alpha$ in a complex neighbourhood of the half-line $(-\infty,Q-\epsilon)$ for each $\epsilon>0$ (see Lemma 3.2). In the next statement, $\beta$ is any number satisfying the above condition.
@@
The expression $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})\in e^{-\beta c}\mathcal{H}$ admits an analytic extension which is regular in a neighbourhood of $\alpha_{1,2}$ and $\alpha_{2,1}$. Moreover,
@@
    $\Phi_{\alpha_{2,1}}(\mathbf{S}_{2Q-\alpha_{2,1}}^{0}\mathds{1})=\left\{\begin{aligned} &\frac{\gamma^{2}}{4}\left(\mu_{\mathrm{L}}^{2}-2\mu_{\mathrm{L}}\mu_{\mathrm{R}}\cos(\pi\frac{\gamma^{2}}{4})+\mu_{\mathrm{R}}^{2}-\mu\sin(\pi\frac{\gamma^{2}}{4})\right)\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}\Psi_{\alpha_{-2,1}}^{\partial}&\text{ if }\gamma<\sqrt{2};\\ &0&\text{if }\gamma>\sqrt{2}.\end{aligned}\right$. (1.5)